# The PanAf-FGBG Dataset: Understanding the Impact of Backgrounds in Wildlife Behaviour Recognition

### v2

@@ 5.3 Background Duration
Experimental Overview. Noting that foreground videos may contain some frames with no animals, effectively supplying background only information for these frames, we aim at quantifying the effect of this on model performance (see Fig. 5). Practically, this is achieved by concatenating increasing numbers of background frames to our foreground videos. To do this, we sampled a set of $\lambda T$ frames from $v^{B}$, where $\lambda$ is a predefined factor controlling the proportion of background frames used. The sampled background frames are denoted as $V^{\mathcal{B}}=\{V_{i_{0}}^{\mathcal{B}},V_{i_{1}}^{\mathcal{B}},\dots,V_{i_{\lambda T-1}}^{\mathcal{B}}\}$, with each index $i_{j}$ selected from $\{0,\dots,T-1\}$. The foreground frames $V^{\mathcal{F}}=\{F_{0}^{\mathcal{F}},F_{1}^{\mathcal{F}},\dots,F_{T-1}^{\mathcal{F}}\}$ were then concatenated with $V^{\mathcal{B}}$, resulting in a combined video sequence $V^{\text{concat}}=\{F_{0}^{\mathcal{F}},F_{1}^{\mathcal{F}},\dots,F_{T-1}^{\mathcal{F}},F_{i_{0}}^{\mathcal{B}},F_{i_{1}}^{\mathcal{B}},\dots,F_{i_{\lambda T-1}}^{\mathcal{B}}\}$ of length $T+\lambda T$. From this concatenated sequence, a starting frame $F_{\text{start}}$ was sampled at random from the indices $\{0,\dots,\lambda T\}$, ensuring that $T$ consecutive frames could be sampled without exceeding the sequence bounds. The final sampled frame sequence $S=\{F_{\text{start}},F_{\text{start}+1},\dots,F_{\text{start}+T-1}\}$ was then used as input for training while retaining the original label $y$ from the foreground video $v^{\mathcal{F}}$.
@@
Analysing the Continuum Between Foreground and Background Videos. Figure 5 shows the full performance comparison between the 2D R50, 3D R50 and MViT-V2 models as background duration increases in videos ($\lambda$). As expected, Figure 5 confirms that as $\lambda$ increases model performance declines towards background model performance (see Tab 1). On $\mathcal{D}^{overlap}$, the 3D-R50 and MViT-V2 models exhibit a sharp decline before beginning to converge. This occurs at $\lambda=32$ and $\lambda=64$ for the 3D-R50 and MViT-V2 models, respectively. Additionally, as shown by the interval between $\lambda=0$ and $\lambda=32$, the uAP and mAP of 3D-R50 falls by 16.88% and 14.32%, respectively, while falling by only 7.23% and 9.13% for the MViT-V2 model. Earlier convergence to near BG-only performance and a more rapid rate of decline, show that 3D-R50 is more sensitive to increases in background durations than MViT-V2. Despite starting from lower performance, the 2D R50 model exhibits a more gradual, near-linear decline, indicating less sensitivity to changes in $\lambda$. While the behaviour of the 2D-R50 and 3D-R50 models remain relatively consistent on $\mathcal{D}^{disjoint}$, the behaviour of MViT-V2 changes significantly. Its uAP declines more steeply, falling from 70.05% to 48.29% (a decrease of 22.21%) between $\lambda=0$ and $\lambda=32$, where its performance converges. It is also more sensitive to smaller increases in background durations, falling from 70.50% to 64.08% (a decrease of 6.42%) between $\lambda=0$ and $\lambda=1.0$, where a decrease of only 0.38% is experienced on $\mathcal{D}^{overlap}$ over the same interval. At higher values of $\lambda$ (i.e., $>40$) the performance drops below that of the 3D-R50 (although only once the background-only performance has been surpassed). Although uAP falls more steeply, mAP declines more slowly and performance above that of the BG-only model is maintained even at the highest values of $\lambda$. This suggests that, although the MViT-V2 achieves higher baseline performance on the disjoint dataset, it is more sensitive to increases in background frames when evaluating uAP – a specific data scenario drowning the advantages of the transformer model in background information.
@@ 5.4 Background Bias Mitigation
Augmenting Performance via Explicit Background Neutralisation. In this final experimental section, we justify the need for paired, real-world background samples, as uniquely provided by our dataset. We show that these sample pairings are demonstrably qualified to significantly improve model performance. To show this, we analyse the impact of background subtraction operations in both input and embedding spaces across each of the model architectures and dataset configurations. Two background-subtracted views of our dataset were created for input space experiments. This was done by performing pixel-wise subtraction of $\mathcal{B}$ and $\tilde{\mathcal{B}}$ from $\mathcal{F}$. To perform algebraic background subtraction in latent space, we extracted the feature vectors of the foreground video pairs. As described in Sec. 5.2, we achieve this by processing the video tensors up to the last convolutional layer and averaging the tokens from the last hidden state in the ResNet and MViT-V2 models, respectively. The background-subtracted embedding is then computed as $z^{\mathcal{F}-\mathcal{B}}=z^{\mathcal{F}}-(1-\alpha)\cdot z^{\mathcal{B}}$ where $\alpha$ is a coefficient used to control the relative magnitude of the algebraic subtraction operation as schematically shown in Fig. 6. In practice, the modulation parameter $\alpha\in[0,1]$ was ablated to increase linearly or exponentially over the course of training and results are reported for both. In all cases $\alpha$ was initialized at zero and increased to one by the end of training.
@@
Table 2: Quantifying Performance Improvements of Background Subtraction. Performance comparison of the 2D R50, 3D R50, and MViT-V2 models on both dataset configurations: $\mathcal{D}_{O}$ and $\mathcal{D}_{D}$. Each model’s performance is reported when training on $\mathcal{F}\ -\mathcal{B}$ or $\mathcal{F}\ -\tilde{\mathcal{B}}$ videos in input (light green) and embedding space (light blue). The highest scores are indicated in bold. Results confirm superiority of the availability of paired, real-world background videos and algebraic operation in embedding space, that is over synthetic data or input space operations.
Results and Performance Improvements. For the basic 2D-R50 model, background subtraction in input space is highly effective. When using $\mathcal{B}$, we observe substantial performance improvements over the baseline, with percentage increases of +8.06% and 5.84% in uAP and mAP on $D^{overlap}$, and +9.06% and +7.76% in uAP and mAP on $D^{disjoint}$. Using $\tilde{\mathcal{B}}$ also yields improvements, though these gains are notably smaller. Background subtraction in embedding space does not provide any benefit to the 2D-R50. Although input space background subtraction is effective for the 2D-R50, it does not improve performance for the 3D-R50 and MViT-V2 models. Neither original nor synthetic backgrounds provide an advantage over their respective baselines, indicating that input space subtraction may not be suitable for architectures that process spatio-temporal features.
@@
6 Limitations
Key limitations include that; (1) our experiments and dataset relate, due to the nature of camera traps, to statically installed sensors only; (2) the latent space background subtraction requires additional data (in the form of background frames) and has computational implications of two forward passes of the model (instead of one) for each training step; (3) the background concatenation experiments may have synthesised videos that are unnatural (e.g., a chimpanzee sitting could suddenly disappear without moving out of frame) and may affect learning performance.
@@ 7 Conclusion
In this work, we introduced PanAf-FGBG – the first wildlife dataset specifically designed to evaluate backgrounds and out-of-distribution generalisation for behaviour recognition. It is a large, richly annotated wild chimpanzee behaviour dataset which defines explicit data configurations for experiments with overlapping and disjoint camera locations. Performing in-depth experimentation with this unique dataset, we demonstrated that real-world background information from wild habitats is not only a powerful predictor of animal behaviour, but can be utilised to significantly improve recognition performance across models.
Beyond these results, the release of this dataset will enable the computer vision community to run direct evaluative studies under in-distribution and out-of-distribution conditions to improve the utilisation of backgrounds for deep action recognition in the wildlife domain. The rich metadata will allow further tasks and experimental configurations to be created. We hope that this will aid model transferability and performance in this challenging domain and ultimately benefit endangered and charismatic species like the one featured in this paper.
@@ 1 Dataset Statistics
We present additional dataset statistics for the country, research site, camera locations and behaviours. Specifically, (i) the distribution of countries and the corresponding research sites are displayed in Fig. 1 (ii) the accumulative proportion of videos contributed by each camera is shown in Fig. 2 and; (iii) a comparison of the behaviour distribution for the overlapping and disjoint datasets is displayed in Fig. 3
@@
Figure 2: Accumulative proportion of videos contributed by each camera.. The y-axis represents the accumulative proportion of videos, with individual cameras arranged by their contribution on the x-axis. The red dashed line divides the cameras into two groups: those contributing the first 50% of the data (left) and the remaining 50% (right). Many videos are contributed by a relatively small number of cameras: 48 cameras contribute 50% of the footage.
Figure 3: Comparison of the proportion of videos containing each behaviour between overlapping and disjoint datasets. Proportions for the overlapping dataset are shown in blue and the disjoint dataset in red. Behaviours are ordered from highest to lowest proportion, with exact values displayed above each bar for easy comparison.
@@ 2 Synthetic Background Generation
We generated synthetic background videos using SAM2 [37] and mean pixel value filling. Specifically, we prompted the SAM2.1-Large model using a single spatial coordinate indicating the location of the chimpanzee to produce an initial segmentation mask. We then leveraged the automatic mask propagation functionality of SAM2 to create spatio-temporal masklets for the full video. Note that spatial coordinates were produced manually. Then, we filled the area indicated by the segmentation mask with the mean pixel value for the frame (see Fig. 4 for examples). Synthetic background videos, spatial coordinates, and segmentation masks will be published with the dataset.

### v3

@@ 5.3 Background Duration
Experimental Overview. Noting that foreground videos may contain some frames with no animals, effectively supplying background only information for these frames, we aim at quantifying the effect of this on model performance (see Fig. 5). Practically, this is achieved by concatenating increasing numbers of background frames to our foreground videos. To do this, we sampled a set of $\lambda T$ frames from $v^{B}$, where $\lambda$ is a predefined factor controlling the proportion of background frames used. The sampled background frames are denoted as $V^{\mathcal{B}}=\{V_{i_{0}}^{\mathcal{B}},V_{i_{1}}^{\mathcal{B}},\dots,V_{i_{\lambda T-1}}^{\mathcal{B}}\}$, with each index $i_{j}$ selected from $\{0,\dots,T-1\}$. The foreground frames $V^{\mathcal{F}}=\{F_{0}^{\mathcal{F}},F_{1}^{\mathcal{F}},\dots,F_{T-1}^{\mathcal{F}}\}$ were then concatenated with $V^{\mathcal{B}}$, resulting in a combined video sequence $V^{\text{concat}}=\{F_{0}^{\mathcal{F}},F_{1}^{\mathcal{F}},\dots,F_{T-1}^{\mathcal{F}},F_{i_{0}}^{\mathcal{B}},F_{i_{1}}^{\mathcal{B}},\dots,F_{i_{\lambda T-1}}^{\mathcal{B}}\}$ of length $T+\lambda T$. From this concatenated sequence, a starting frame $F_{\text{start}}$ was sampled at random from the indices $\{0,\dots,\lambda T\}$, ensuring that $T$ consecutive frames could be sampled without exceeding the sequence bounds. The final sampled frame sequence $S=\{F_{\text{start}},F_{\text{start}+1},\dots,F_{\text{start}+T-1}\}$ was then used as input for training while retaining the original label $y$ from the foreground video $v^{\mathcal{F}}$. Background frames are appended because motion-triggered cameras are more likely to capture activity first, resulting in a more natural ordering.
Figure 5: Effect of Increasing Background Duration on Performance. Comparison of 2D R50, 3D R50 and MViT-V2 model performance as a function of the simulated background duration ($\lambda$). The 3D-R50 is more sensitive to increasing background durations than MViT-V2. MViT-V2 is relatively robust to increasing background durations on the overlapping dataset, although it becomes highly sensitive when evaluated on OOD data.
@@
Analysing the Continuum Between Foreground and Background Videos. Fig. 5 shows the performance comparison between the 2D R50, 3D R50 and MViT-V2 models as background duration increases in videos ($\lambda$). As expected, Fig. 5 confirms that as $\lambda$ increases model performance declines towards background model performance (see Tab 1). On $\mathcal{D}^{overlap}$, the 3D-R50 and MViT-V2 models exhibit a sharp decline before beginning to converge. This occurs at $\lambda=32$ and $\lambda=64$ for the 3D-R50 and MViT-V2 models, respectively. Additionally, as shown by the interval between $\lambda=0$ and $\lambda=32$, the uAP and mAP of 3D-R50 falls by 16.88% and 14.32%, respectively, while falling by only 7.23% and 9.13% for the MViT-V2 model. Earlier convergence to near BG-only performance and a more rapid rate of decline, show that 3D-R50 is more sensitive to increases in background durations than MViT-V2. Despite starting from lower performance, the 2D R50 model exhibits a more gradual, near-linear decline, indicating less sensitivity to changes in $\lambda$. While the behaviour of the 2D-R50 and 3D-R50 models remain relatively consistent on $\mathcal{D}^{disjoint}$, the behaviour of MViT-V2 changes significantly. Its uAP declines more steeply, falling from 70.05% to 48.29% (a decrease of 22.21%) between $\lambda=0$ and $\lambda=32$, where its performance converges. It is also more sensitive to smaller increases in background durations, falling from 70.50% to 64.08% (a decrease of 6.42%) between $\lambda=0$ and $\lambda=1.0$, where a decrease of only 0.38% is experienced on $\mathcal{D}^{overlap}$ over the same interval. At higher values of $\lambda$ (i.e., $>40$) the performance drops below that of the 3D-R50 (although only once the background-only performance has been surpassed). Although uAP falls more steeply, mAP declines more slowly and performance above that of the BG-only model is maintained even at the highest values of $\lambda$. This suggests that, although the MViT-V2 achieves higher baseline performance on the disjoint dataset, it is more sensitive to increases in background frames when evaluating uAP; a scenario where the model’s advantage is drowned out by background information.
@@ 5.4 Background Bias Mitigation
Augmenting Performance via Explicit Background Neutralisation. In this final experimental section, we justify the need for paired, real-world background samples, as uniquely provided by our dataset. We show that these sample pairings are able to significantly improve model performance. To show this, we analyse the impact of background subtraction operations in both input and embedding spaces across each of the model architectures and dataset configurations. Two background-subtracted views of our dataset were created for input space experiments. This was done by performing pixel-wise subtraction of $\mathcal{B}$ and $\tilde{\mathcal{B}}$ from $\mathcal{F}$. To perform algebraic background subtraction in latent space, we extracted the feature vectors of the foreground video pairs. As described in Sec. 5.2, we achieve this by processing the video tensors up to the last convolutional layer and averaging the tokens from the last hidden state in the ResNet and MViT-V2 models, respectively. The background-subtracted embedding is then computed as $z^{\mathcal{F}-\mathcal{B}}=z^{\mathcal{F}}-(1-\alpha)\cdot z^{\mathcal{B}}$ where $\alpha$ is a coefficient used to control the relative magnitude of the algebraic subtraction operation as schematically shown in Fig. 6. In practice, the modulation parameter $\alpha\in[0,1]$ was ablated to increase linearly or exponentially over the course of training and results are reported for both. In all cases $\alpha$ was initialized at zero and increased to one by the end of training.
@@
Table 2: Quantifying Performance Improvements of Background Subtraction. Comparison of the 2D R50, 3D R50, and MViT-V2 models on both dataset configurations. Performance is reported when training on $\mathcal{F}\ -\mathcal{B}$ or $\mathcal{F}\ -\tilde{\mathcal{B}}$ videos in input (light green) and embedding space (light blue). First and second highest scores are indicated for each model, while the highest overall scores are marked by an asterisks (*). Note the baseline is trained on $\mathcal{F}$, as reported in Tab. 1. $\mathcal{F}+\mathcal{B}$ provides an alternative baseline but is reliant on background videos at test time. Results confirm the superiority of real-world background videos over synthetic data, and of latent space neutralisation over input space background subtraction.
Results and Performance Improvements. Tab. 2 shows that for the basic 2D-R50 model, background subtraction in input space is highly effective. When using $\mathcal{B}$, we observe substantial performance improvements over the baseline, with percentage increases of +8.06% and 5.84% in uAP and mAP on $D^{overlap}$, and +9.06% and +7.76% in uAP and mAP on $D^{disjoint}$. Using $\tilde{\mathcal{B}}$ also yields improvements, though these gains are notably smaller. Background subtraction in embedding space does not provide any benefit to the 2D-R50. Although input space background subtraction is effective for the 2D-R50, it does not improve performance for the 3D-R50 and MViT-V2 models. Neither original nor synthetic backgrounds provide an advantage over their respective baselines, indicating that input space subtraction may not suit spatio-temporal architectures.
@@ 6 Conclusion
In this work, we introduced PanAf-FGBG, the first wildlife dataset designed to assess backgrounds and out-of-distribution generalization in behavior recognition. Our results demonstrate that real-world background information is a strong predictor of animal behavior and can significantly enhance recognition performance. While there are some limitations — such as its restriction to statically installed camera trap videos, the additional data and compute required for latent space background subtraction, and the potential for unnatural artifacts in background concatenation — the dataset remains a valuable resource. PanAf-FGBG enables direct evaluation under out-of-distribution conditions, and its rich metadata supports additional tasks and experimental setups. We hope it will help with understanding and improving model performance in this challenging domain and ultimately benefit endangered and charismatic species like the one featured in this paper.
Acknowledgements
We thank the Pan African Programme: ‘The Cultured Chimpanzee’ team and its collaborators for allowing the use of their data for this paper. We thank Amelie Pettrich, Antonio Buzharevski, Eva Martinez Garcia, Ivana Kirchmair, Sebastian Schütte, Linda Gerlach and Fabina Haas. We also thank management and support staff across all sites; specifically Yasmin Moebius, Geoffrey Muhanguzi, Martha Robbins, Henk Eshuis, Sergio Marrocoli and John Hart. Thanks to the team at https://www.chimpandsee.org particularly Briana Harder, Anja Landsmann, Laura K. Lynn, Zuzana Macháčková, Heidi Pfund, Kristeena Sigler and Jane Widness. The work that allowed for the collection of the dataset was funded by the Max Planck Society, Max Planck Society Innovation Fund, and Heinz L. Krekeler. In this respect we would like to thank: Ministre des Eaux et Forêts, Ministère de l’Enseignement supérieur et de la Recherche scientifique in Côte d’Ivoire; Institut Congolais pour la Conservation de la Nature, Ministère de la Recherche Scientifique in Democratic Republic of Congo; Forestry Development Authority in Liberia; Direction Des Eaux Et Forêts, Chasses Et Conservation Des Sols in Senegal; Makerere University Biological Field Station, Uganda National Council for Science and Technology, Uganda Wildlife Authority, National Forestry Authority in Uganda; National Institute for Forestry Development and Protected Area Management, Ministry of Agriculture and Forests, Ministry of Fisheries and Environment in Equatorial Guinea. This work was supported by the UKRI CDT in Interactive AI (grant EP/S022937/1). This work was in part supported by the US National Science Foundation Awards No. 2118240 "HDR Institute: Imageomics: A New Frontier of Biological Information Powered by Knowledge-Guided Machine Learning" and Award No. 2330423 and Natural Sciences and Engineering Research Council of Canada under Award No. 585136 for the “AI and Biodiversity Change (ABC) Global Center”.
@@ 1 Dataset Statistics
We present additional dataset statistics for the country, research site, camera locations and behaviours. Specifically, (i) the distribution of countries and the corresponding research sites are displayed in Fig. 1; (ii) a comparison of the behaviour distribution for the overlapping and disjoint datasets is displayed in Fig. 2; (iii) time interval standardisation statistics are shown in Fig 3 and; (iv) the accumulative proportion of videos contributed by each camera is shown in Fig. 4.
@@
Figure 2: Comparison of the proportion of videos containing each behaviour between overlapping and disjoint datasets. Behaviours are ordered from highest to lowest proportion, with exact values displayed above each bar for easy comparison.
Figure 3: Time interval standardisation between foreground and background video pairs. The majority of foreground background video pairs are sampled within one hour of each other.
Figure 4: Accumulative proportion of videos contributed by each camera.. The y-axis represents the accumulative proportion of videos, with individual cameras arranged by their contribution on the x-axis. The red dashed line divides the cameras into two groups: those contributing the first 50% of the data (left) and the remaining 50% (right).
@@ 2 Synthetic Background Generation
We generated synthetic background videos using SAM2 [44] and mean pixel value filling. Specifically, we prompted the SAM2.1-Large model using a single spatial coordinate indicating the location of the chimpanzee to produce an initial segmentation mask. We then leveraged the automatic mask propagation functionality of SAM2 to create spatio-temporal masklets for the full video. Note that spatial coordinates were produced manually. Then, we filled the area indicated by the segmentation mask with the mean pixel value for the frame (see Fig. 5 for examples).